# Joint AP-UE Association and Power Factor Optimization for Distributed Massive MIMO

### v2

@@ II System Model
[…]
@@ II-A Uplink Pilot Training
[…]
$\textbf{y}^{pilot}_{m}=\sum_{t\in\mathcal{T}}\textbf{g}_{mt}\mathbf{\Psi}_{t}+\textbf{n}_{m},$
[…]
$\hat{\textbf{g}}_{mt}=\frac{\beta_{mk}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\textbf{y}^{pilot}_{m}\mathbf{\Psi}_{t}.$
@@
$\gamma_{mt}=\left(\frac{p_{p}\tau\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}p_{p}\tau\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\right).$
@@ II-C Problem Formulation
where $\bm{\eta^{u}}=[\eta^{u}_{1},\eta^{u}_{2},...\eta^{u}_{T}]^{\intercal}$ and D is the AP-UE association matrix with each element as $d_{mt}$. $\textbf{D}_{t}$ represents the column vector of matrix of D with respect to the UE $t$. The regularization coefficient $\alpha\geq 0$ is the balancing factor between the front-haul load and the SE. $\text{SE}^{\text{QoS}}_{t}$ represents the minimum QoS requirement for the UE $t$. The constraint (4b) represents the binary variable indicating the association between the UEs and APs. Constraint (4c) represents the uplink power control coefficient. Constraint (4d) is for the minimum QoS requirement and constraint (4e) specifies that atleast one AP must be connected to the UE. The objective function in (4a) is a non-smooth and mixed integer type non-convex function that makes the overall problem NP-hard. Finding the optimal solution of such mixed integer type problems requires high computation resources. In the following section we provide a heuristic solution of problem (II-C) using a relaxation approach.
@@ III Proposed Joint Solution using Relaxation Technique
Problem (II-C) is a mixed-integer type problem which has a very high complexity, so we first convert the discrete binary constraint (4b) into a continuous one, resulting in:
@@
After computing the stationary point of (8) with respect to $\mathbf{\Gamma}$, we get the optimal value of $\lambda^{*}_{t}=\frac{w^{\prime}}{1+{\Gamma^{*}_{t}}}$. where $w^{\prime}=w/{\ln 2}$ with the optimal value of $\mathbf{\Gamma}=\mathbf{\Gamma^{*}}$, where $\mathbf{\Gamma^{*}}=[\Gamma^{*}_{1},\Gamma^{*}_{2},...\Gamma^{*}_{T}]^{\intercal}$. Substituting this $\lambda^{*}_{t}$ in (8) and using the Lagrangian dual transformation [17], the problem in (III) can be reformulated as :
@@
It can be observed that after substituting the optimal value of $\mathbf{\Gamma}$, as in (2), in equation (9a), the optimization problem in (III) reduces exactly to the one in (III). If we look at optimization problem (III), we still have some non-linearities and non-convexities due to a ratio term and a product of two variables in the objective function. First, we attempt to handle the ratio term in the objective function using the quadratic transformation approach [17]. The quadratic transformation is performed on the objective function (9a) such that it becomes:
@@
After substituting the optimal value of u in equation (10) and then solving optimization problem (III) is equivalent to solving Problem (III). Notice that Problem (III) is a convex optimization problem in different blocks. If we fix D, $\mathbf{\Gamma}$ and u, then the objective function of problem (12a) becomes concave with respect to $\bm{\eta^{u}}$. Then, the problem reduces to:
@@
Similarly, upon fixing $\bm{\eta^{u}}$, $\mathbf{\Gamma}$ and u, the objective function of (III) becomes concave with respect to D. Then, the optimization problem becomes:
@@
Problem III can be solved following the proposed iterative approach as follows. This problem is decomposed into two distinct blocks, denoted as Problem III and Problem III. Initially, the solution for Problem III is obtained while keeping D, $\mathbf{\Gamma}$, and u fixed. Subsequently, based on the solution derived for Problem III, the solution for Problem III is determined. This is repeated until the convergence criterion is met. Algorithm 1 provides the details of this approach.
1:  Feasible Initialization: ${\bm{\eta_{k}^{u}}}^{(0)}$, $\textbf{D}^{(0)}$ and $i=0$, $\forall k$.
@@
7:     Solve optimization problem in (III) for ${\bm{\eta_{k}^{u}}}^{(i)}$.
@@
11:     Solve optimization problem in (III) for $\textbf{D}^{(i)}$.
@@ V Conclusion
Convergence Analysis : For the convergence analysis of the Algorithm 1, the primary goal is to prove that both optimization problem defined in (III) and (III) converges to the their respective optimal solutions. Thereafter, we prove that in each iteration of Algorithm 1, the optimization function in (12a) is non-decreasing with respect to $\mathbf{\eta^{u}}$ and D.

### v3

@@ II System Model
We consider a system consisting of $T$ single antenna UEs and $M$ multi-antenna APs, where each AP has $A$ antennas such that $T<<MA$. The APs and UEs are distributed uniformly over the given coverage area and APs are jointly serving all users over the same time and frequency resources. In order to help in the signal processing and coordination among APs, they are connected with a CPU via a front-haul connection. We consider the TDD mode of operation where the AP-UE channel vector is estimated using the uplink transmission. To model our channel, we have considered the block fading model with a coherence block of length $L_{c}$. Let $L_{p}$ be the length of orthogonal pilots such that $L_{p}$ length of coherence block is used for uplink pilot training. The Rayleigh fading coefficient representing the small-scale fading is denoted by $\textbf{h}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$, whereas the large scale-fading coefficient (LSFC) that incorporates both path-loss and shadowing effects is denoted by $\beta_{mt}$. The channel vector $\textbf{g}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$, that incorporates both the small-scale fading as well as large scale fading, is defined as $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$. We assume that $\textbf{h}_{t,m}{\sim}\mathcal{N}_{\mathbb{C}}(0,1)$ is independent and identically distributed (i.i.d.) random vector.
@@ II-A Uplink Pilot Training
We assume that during uplink channel estimation, each UE $t\in\mathcal{T}\overset{\Delta}{=}\{1,...,T\}$ transmits the pilot sequence $\sqrt{p_{p}L_{p}\eta_{t}}\mathbf{\Psi}_{t}\in\mathbb{C}^{L_{p}\times 1}$, such that $||\mathbf{\Psi}_{t}||^{2}=1$, where $p_{p}$ represents the normalized signal-to-noise power. The received signal $\textbf{y}^{pilot}_{m}\in\mathbb{C}^{\textit{A}\times L_{p}}$ at the $m^{th}$ AP is given by :
$\textbf{y}^{pilot}_{m}=\sum_{t\in\mathcal{T}}\sqrt{p_{p}L_{p}\eta_{t}}\textbf{g}_{mt}\mathbf{\Psi}_{t}+\textbf{n}_{m},$
where $\textbf{n}_{m}\in\mathbb{C}^{\textit{A}\times L_{p}}$ is the complex additive white Gaussian noise matrix of the AP $m$ consists of complex i.i.d. elements. The MMSE estimate $\hat{\textbf{g}}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$ of $\textbf{g}_{mt}$ is given by [3]:
$\hat{\textbf{g}}_{mt}=\frac{\sqrt{p_{p}L_{p}\eta_{t}}\beta_{mk}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\eta_{t}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\textbf{y}^{pilot}_{m}\mathbf{\Psi}_{t}.$
@@
$\gamma_{mt}=\left(\frac{p_{p}L_{p}\eta_{t}\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\eta_{t}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\right).$
@@ II-C Problem Formulation
where $\bm{\eta^{u}}=[\eta^{u}_{1},\eta^{u}_{2},...\eta^{u}_{T}]^{\intercal}$ and D is the AP-UE association matrix with each element as $d_{mt}$. $\textbf{D}_{t}$ represents the column vector of matrix of D with respect to the UE $t$. The regularization coefficient $\alpha\geq 0$ is the balancing factor between the front-haul load and the SE. $\text{SE}^{\text{QoS}}_{t}$ represents the minimum QoS requirement for the UE $t$. The constraint (4b) represents the binary variable indicating the association between the UEs and APs. Constraint (4c) represents the uplink power control coefficient. Constraint (4d) is for the minimum QoS requirement and constraint (4e) specifies that atleast one AP must be connected to the UE. The objective function in (4a) is a non-smooth and mixed integer type non-convex function that makes the overall problem NP-hard. Finding the optimal solution of such mixed integer type problems requires high computation resources. In the following section we provide a heuristic solution of problem (4) using a relaxation approach.
@@ III Proposed Joint Solution using Relaxation Technique
Problem (4) is a mixed-integer type problem which has a very high complexity, so we first convert the discrete binary constraint (4b) into a continuous one, resulting in:
@@
After computing the stationary point of (8) with respect to $\mathbf{\Gamma}$, we get the optimal value of $\lambda^{*}_{t}=\frac{w^{\prime}}{1+{\Gamma^{*}_{t}}}$. where $w^{\prime}=w/{\ln 2}$ with the optimal value of $\mathbf{\Gamma}=\mathbf{\Gamma^{*}}$, where $\mathbf{\Gamma^{*}}=[\Gamma^{*}_{1},\Gamma^{*}_{2},...\Gamma^{*}_{T}]^{\intercal}$. Substituting this $\lambda^{*}_{t}$ in (8) and using the Lagrangian dual transformation [17], the problem in (6) can be reformulated as :
@@
It can be observed that after substituting the optimal value of $\mathbf{\Gamma}$, as in (2), in equation (9a), the optimization problem in (9) reduces exactly to the one in (6). If we look at optimization problem (9), we still have some non-linearities and non-convexities due to a ratio term and a product of two variables in the objective function. First, we attempt to handle the ratio term in the objective function using the quadratic transformation approach [17]. The quadratic transformation is performed on the objective function (9a) such that it becomes:
@@
After substituting the optimal value of u in equation (10) and then solving optimization problem (12) is equivalent to solving Problem (9). Notice that Problem (12) is a convex optimization problem in different blocks. If we fix D, $\mathbf{\Gamma}$ and u, then the objective function of problem (12a) becomes concave with respect to $\bm{\eta^{u}}$. Then, the problem reduces to:
@@
Similarly, upon fixing $\bm{\eta^{u}}$, $\mathbf{\Gamma}$ and u, the objective function of (12) becomes concave with respect to D. Then, the optimization problem becomes:
@@
Problem 12 can be solved following the proposed iterative approach as follows. This problem is decomposed into two distinct blocks, denoted as Problem 13 and Problem 14. Initially, the solution for Problem 13 is obtained while keeping D, $\mathbf{\Gamma}$, and u fixed. Subsequently, based on the solution derived for Problem 13, the solution for Problem 14 is determined. This is repeated until the convergence criterion is met. Algorithm 1 provides the details of this approach.
1:  Feasible Initialization: ${\bm{\eta_{t}^{u}}}^{(0)}$, $\textbf{D}^{(0)}$ and $i=0$, $\forall t$.
@@
7:     Solve optimization problem in (13) for ${\bm{\eta_{k}^{u}}}^{(i)}$.
@@
11:     Solve optimization problem in (14) for $\textbf{D}^{(i)}$.
@@ V Conclusion
Convergence Analysis : For the convergence analysis of the Algorithm 1, the primary goal is to prove that both optimization problem defined in (13) and (14) converges to the their respective optimal solutions. Thereafter, we prove that in each iteration of Algorithm 1, the optimization function in (12a) is non-decreasing with respect to $\mathbf{\eta^{u}}$ and D.